OpenStudy (anonymous):

use the change of base formula to evaluate log3(33)

OpenStudy (zehanz):

So you can get log based 10 on your calculator?

OpenStudy (anonymous):

\[\log_{3}(33)=\frac{ \log(33) }{ \log(3) } \approx 3.183 \]

OpenStudy (zehanz):

The formula to change base ... has been given by@genius12. Unfortunately, he also gave away the answer...